ganeshie8 (ganeshie8):

\[\sum\limits_{d\mid n}\mu^2(d)/\tau(d)=?\]

oregonduck (oregonduck):

\(\color{blue}{\text{Originally Posted by}}\) @ganeshie8 \[\sum\limits_{d\mid n}\mu^2(d)/\tau(d)=?\] \(\color{blue}{\text{End of Quote}}\) It is blank for me

ganeshie8 (ganeshie8):

The functions \(\mu\) and \(\tau\) are multiplicative, it follows that \(\mu^2/\tau\) is also multiplicative.

ganeshie8 (ganeshie8):

If \(n = p_1^{e_1}p_2^{e_2}\cdots p_r^{e_r}\), then we have : \[\sum\limits_{d\mid n}\mu^2(d)/\tau(d)=\left(\sum\limits_{d\mid p_1^{e_1}}\mu^2(d)/\tau(d)\right)\left(\sum\limits_{d\mid p_2^{e_2}}\mu^2(d)/\tau(d)\right)\cdots\left(\sum\limits_{d\mid p_r^{e_r}}\mu^2(d)/\tau(d)\right) \]

ganeshie8 (ganeshie8):

thanks @OregonDuck

imqwerty (imqwerty):

we need to prove LHS=RHS?

ganeshie8 (ganeshie8):

we need to find an expression for the sum in main question

ganeshie8 (ganeshie8):

the sum itself is multiplicative, meaning : \[f(ab) = f(a)f(b)\] whenever \(\gcd(a,b)=1\)

OpenStudy (pawanyadav):

Here is no LHS and RHS

ganeshie8 (ganeshie8):

since all the prime powers are relatively prime, we can split the sum as shown above..

OpenStudy (pawanyadav):

Don't understand what it is? The previous one is at least understandable but this one is so far..............

imqwerty (imqwerty):

:) we can write it in this form-\(\sum\limits_{d\mid n}\mu^2(d)/\tau(d)=\left(\sum\limits_{d\mid p_1^{e_1}}\mu^2(d)/\tau(d)\right)\left(\sum\limits_{d\mid p_2^{e_2}}\mu^2(d)/\tau(d)\right)\cdots\left(\sum\limits_{d\mid p_r^{e_r}}\mu^2(d)/\tau(d)\right)\) the divisors of any \(p_i^{e_i}\) will be like this-> \(1,~p_i,~p_i^2,~p_i^3,..\) \(\left(\sum\limits_{d\mid p_1^{e_1}}\mu^2(d)/\tau(d)\right)\) = \(\Large{\frac{ \mu^2(1) }{ \tau(1) }+\frac{ \mu^2(p_1) }{ \tau(p_1) }+\frac{ \mu^2(p_1^2) }{ \tau(p_1^2) }}+\frac{ \mu^2(p_1^3) }{ \tau(p_1^3) }....\) after \(p_1^2,~p_1^3,~p_1^4.....\) are divisible by \(p^2\) so \(0=\mu(p_1^2)=\mu(p_1^3)=\mu(p_1^4)=...\) so basically all that part will become 0 and we will be left with \(\left(\sum\limits_{d\mid p_1^{e_1}}\mu^2(d)/\tau(d)\right)=\Large{\frac{ \mu^2(1) }{ \tau(1) }+\frac{ \mu^2(p_1) }{ \tau(p_1) }}\) \(\left(\sum\limits_{d\mid p_1^{e_1}}\mu^2(d)/\tau(d)\right)=\large{1+\frac{1}{2}}\) if the prime factorization of \(n\) has \(r\) primes then \(\sum\limits_{d\mid n}\mu^2(d)/\tau(d)=\large{\left( 1+\frac{ 1 }{ 2 } \right)^r}\) we had to do this right?

ganeshie8 (ganeshie8):

wow! that looks perfect !

ganeshie8 (ganeshie8):

but \(r\) is a variable that is not there in the actual question right ?

ganeshie8 (ganeshie8):

that can be easily fixed by using number of distinct prime factors function : \(\omega(n)\) http://mathworld.wolfram.com/DistinctPrimeFactors.html

ganeshie8 (ganeshie8):

\[\sum\limits_{d\mid n}\mu^2(d)/\tau(d)=\large{\left( 1+\frac{ 1 }{ 2 } \right)^{\omega(n)}}\]

ganeshie8 (ganeshie8):

\(\omega(n)\) spits out the number of distinct prime factors of \(n\)

ganeshie8 (ganeshie8):

\(\omega(2*5*7^5) = 3\) \(\omega(11*31) = 2\) etc

imqwerty (imqwerty):

ok so instead of \(r\) we should use the function \(\omega(n)\) because its a general type of notation with our variable \(n\)

ganeshie8 (ganeshie8):

its not a big deal, \(\omega(n)\) is a very popular function, it shows up in many places

imqwerty (imqwerty):

okay :) i'll study this function

ganeshie8 (ganeshie8):

not much to study actually, it just spits out the number of prime factors that all

ganeshie8 (ganeshie8):

\(\tau, \sigma, \phi,\mu\) are more interesting functions...

ganeshie8 (ganeshie8):

thats just in my opinion again..

imqwerty (imqwerty):

but what if the number is very big is there any formula to find the number of prime factors

imqwerty (imqwerty):

without doing any prime factorization..

ganeshie8 (ganeshie8):

do you expect there exists such a formula ?

ganeshie8 (ganeshie8):

If such a formula had existed, prime numbers wouldn't be any fun

ganeshie8 (ganeshie8):

because then it can be used to check if a number is prime

imqwerty (imqwerty):

ah yes :)

ganeshie8 (ganeshie8):

people have been trying for such a formula though

ganeshie8 (ganeshie8):

it seems impossible as prime numbers are just so mysterious

ganeshie8 (ganeshie8):

it has this trivial yet interesting property : \[\omega(ab) = \omega(a)+\omega(b)\] whenever \(\gcd(a,b)=1\)

imqwerty (imqwerty):

yeah if gcd is 1 means no common prime factor if prime factorization is like this- \(a=p_1p_2p_3..~~and~~~b=p_ap_bp_c..\) its clear that in the prime factorization of \(ab\) will have all these diff primes so \(\omega(ab)=\omega(a)+\omega(b)\)

OpenStudy (ikram002p):

i can conclude these stuff only :- \(\Large if~ n=p_1p_2....p_i \) such that \( {p_i}'s \) are distinct primes, then \(\Large \sum_{d|n}\dfrac{\mu^2(d)}{\tau(d)}=\sum_{d|n}\dfrac{1}{\tau(d)}\\\Large =\sum_{d|n}\frac{1}{\log_d{\prod_{i|d}{i}}}\) and this also goes for any n. \(\Large \sum_{d|n}\dfrac{\mu^2(d)}{\tau(d)}=\Large \sum_{d|n^2}\dfrac{\mu^2(d)}{\tau(d)}\)

OpenStudy (ikram002p):

wait its like this function be simplified as sum of 1 over tau(d) for d|n, as \(\mu(d)=0 ~~ when ~~p^2|d.\)

OpenStudy (ikram002p):

ok here is a full proof for what i'm saying "need time to write"

OpenStudy (ikram002p):

\( \Large \mu^2(d)=\left\{\begin{matrix} 0 & if & p_i^2|d\\ 1 & else& \end{matrix}\right. \\ \Large \begin{matrix} n=\prod_{i=1}^{\omega(n)}p_i^{k_i} \end{matrix}\\ \Large \begin{matrix} n_{\omega(n)}=\prod_{i=1}^{\omega(n)}p_i \end{matrix}\\ \begin{align*} f(n) &= \sum_{d|n}\frac{\mu^2(d)}{\tau(d)} \\ &= \frac{\mu^2(p_1)}{\tau(p_1)}+\frac{\mu^2(p_1^2)}{\tau(p_1^2)}+...+ \frac{\mu^2(p_1^{k_{\omega(n)} } )}{\tau(p_1^{k_{\omega(n)} } )}\\ &+\frac{\mu^2(p_1p_2)}{\tau(p_1p_2)}+\frac{\mu^2(p_1^2p_2)}{\tau(p_1^2p_2)}+...+ \frac{\mu^2(p_1^{k_{\omega(n)} } p_2^{k_{\omega(n)} } )}{\tau(p_1^{k_{\omega(n)} } p_2^{k_{\omega(n)} } )} \\ &+... \\ &+ \frac{\mu^2(p_1p_2...p_{\omega(n)})}{\tau(p_1p_2...p_{\omega(n)})}+...+\frac{\mu^2((p_1p_2...p_{\omega(n)})^{\omega(n)})}{\tau((p_1p_2...p_{\omega(n)})^{\omega(n)})}\\ &= \frac{1}{\tau(p_1)}+ \frac{1}{\tau(p_2)}+...+\frac{1}{\tau(p_{\omega(n)})} \\ &+\frac{1}{\tau(p_1p_2)}+ \frac{1}{\tau(p_2p_3)} +... \\ &+... \\ &+ \frac{1}{\tau(p_1p_2...p_{\omega(n)})}\\ &=\sum_{d|{n_{\omega(n)}}}\frac{1}{\tau(d)} \end{align*} \)

OpenStudy (ikram002p):

=)

OpenStudy (kainui):

Small note before I really play round with this, is @ganeshie8 's thing, I noticed a while back that the similar \(\Omega(n)\) function is a lot like the logarithm: \[\Omega(p^aq^b)=a \Omega(p) + b \Omega(q)\] Anyways time to play around with the actual problem hehe...!

OpenStudy (ikram002p):

but i have consider about @imqwerty everything looks fine then i thought of such cases like n=1*2*3*5 cuz i have answer of (1+1/2^5+1/2^6+1/2^3) seems contradict with urs if its (1+1/2)3 hmmm

OpenStudy (kainui):

The way I remember it is like this: \[\omega(n) \le \Omega(n)\] the little one is smaller than the bigger one haha. the smaller one strips off the exponents and just counts the unique prime factors while the big one tells you the total prime factors. Pretty cool.

OpenStudy (ikram002p):

oh yeah i should remember that in its easyxD Omega vs omega

imqwerty (imqwerty):

hmm my method is also not applicable for the numbers of the form \(n=p^k\) where \(p\) is a prime. For example n=8

OpenStudy (kainui):

p=2 and k=3 does it not? #_#

imqwerty (imqwerty):

oops sorry :)

OpenStudy (kainui):

Heh :P

OpenStudy (ikram002p):

well my method says if n=p^k and p is prime then f(n)=1+1/2 always xD so i think its the same as urs lol

OpenStudy (ikram002p):

now lets play with multiplicative thingy =D

OpenStudy (kainui):

Oh here's a fun thing for @imqwerty to prove for multiplicative functions if he hasn't seen this before: All interesting multiplicative functions have f(1)=1. Prove it! :P

OpenStudy (ikram002p):

f(mn)=f(n)*f(m) g(m)=f(m*1)=f(m)*f(1)

OpenStudy (ikram002p):

now\( f(n)=\sum_{d|n}1/\tau(n)\) is multiplicative right ?

OpenStudy (kainui):

BTW Here's how I derived the formula for the thing earlier, first I realize once you figure out how a multiplicative function works for \(f(p^k)\) then you know how it works for any combination cause you can multiply them together (I really think in terms of linear algebra and how a transformation acts on the basis set tells you how it affects any arbitrary value here, cause I'm sick in the head like that) So now we use the fact that \(\mu(p^k) = 0\) for k>1 and really all we have to look at are: \[\frac{\mu^2(1)}{\tau(1)}+ \frac{\mu^2(p)}{\tau(p)} = \frac{3}{2}\] Now like I said I do earlier, I multiply the result of acting on one prime altogether, and since there is one of these for each distinct prime: \[\left( \frac{3}{2} \right)^{\omega(n)}\]

imqwerty (imqwerty):

:O ok all interesting functions are-> \(\tau(n)\), \(\phi(n)\), \(\sigma(n)\), \(\mu(n)\) ok so the function is mutiplicative hmmm so (; ikram gave a hint well its pretty clear \(f(n)=f(n*1)=f(n)*f(1)\) can we just say that for f(n) to be f(n) , f(1) must be 1 :/ or else if its not 1 then things get bad for example \(2=\tau(3)=\tau(3*1)=\tau(3)* \tau(1)=2* \tau(1)\) if \(\tau(1)\) is not 1 then \(\tau(3) \ne \tau(3)\) so \(f(1)=1\)

OpenStudy (kainui):

@imqwerty Yeah that's multiplicative. In fact, all multiplicative functions with the Dirichlet convolution are multiplicative functions. What's the Dirichlet convolution? Well one way to write it (in a fairly unintuitive way is): \[h(n) = \sum_{d|n} f(d) g(\tfrac{n}{d}) \] So in your example you can pick \[f(n)=\frac{1}{\tau(n)}\] and \[g(n)=1\] and because of this rule (I have given with no justification :P ) that means h(n) is also multiplicative!

OpenStudy (kainui):

Ahhh @imqwerty Actually although those are some interesting functions, they're not all haha. I made it as a joke based on how I proved it: \[f(1)=f(1*1)=f(1)*f(1)\] So now either f(1)=1 or f(1)=0. But f(1)=0 means f(n)=0 for all values! BORING! xD

OpenStudy (ikram002p):

\( f(n)=f(p_1p_2p_3...p_{\omega({n})}) =f(p_1)f(p_2)...f(p_{\omega(n)})=(1+1/2)^{\omega(n)} \)

OpenStudy (ikram002p):

same as @imqwerty hmm im kinda not convinced about it lol

imqwerty (imqwerty):

@kainui but how can we if f(1) will either be 1 or 0 it wuld depend on the function

OpenStudy (ikram002p):

hahaha there is no multiplicative function with f(1)=0 unless ur working on a base such that f(i)=i identity and i is not 1 .

OpenStudy (kainui):

\[f(1)=f(1*1)\] There I haven't done anything special, 1=1*1 \[f(1*1)=f(1)*f(1)\] since \(\gcd(1,1)=1\) we can split it because f(n) is multiplicative! Now we know \[f(1)=f(1)*f(1)\] and we can divide \(f(1)\) from both sides to see it must be 1 as long as f(1) is no 0. If f(1)=0 then we have for all values: \[f(n)=f(1*n)=f(1)*f(n)=0*f(n)=0\]

imqwerty (imqwerty):

:O okay :D

imqwerty (imqwerty):

what culd be the possible reason behind the failure of (1+1/2)^r method?

OpenStudy (ikram002p):

well, two reasons major error misunderstood

OpenStudy (kainui):

I don't think there is a failure of the method, maybe this could help convince you that we can do this thing where we separate it and then combine it back together: Multiply the left side to show that you get the right hand side for this arbitrary multiplicative function: \[[f(1)+f(p)]*[f(1)+f(q)+f(q^2)]=\]\[ f(1)+f(p)+f(q)+f(pq)+f(q^2)+f(pq^2) \]

imqwerty (imqwerty):

i think its not multiplicative \(\sum\limits_{d\mid n}\mu^2(d)/\tau(d)=\left(\sum\limits_{d\mid p_1^{e_1}}\mu^2(d)/\tau(d)\right)\left(\sum\limits_{d\mid p_2^{e_2}}\mu^2(d)/\tau(d)\right)\cdots\left(\sum\limits_{d\mid p_r^{e_r}}\mu^2(d)/\tau(d)\right)\) i plugged n=30 in here and got the same answer as \(\large\ (1+\frac{1}{2})^3\)

OpenStudy (kainui):

You think it's not multiplicative but you got the same answer...?

OpenStudy (ikram002p):

f(1*2*3)= 1+1/2+1/2+1/4= 9/4 multiplicative method says f(1*2*3)=(3/2)^2=9/8 hahahahaha ok we have calculation error only :P

imqwerty (imqwerty):

i mean like i put n=30 in that 1st express and got answer 27/64 but it should be 13/4 so something is wrong there

OpenStudy (ikram002p):

which expression ?

imqwerty (imqwerty):

this one- \(\sum\limits_{d\mid n}\mu^2(d)/\tau(d)=\left(\sum\limits_{d\mid p_1^{e_1}}\mu^2(d)/\tau(d)\right)\left(\sum\limits_{d\mid p_2^{e_2}}\mu^2(d)/\tau(d)\right)\cdots\left(\sum\limits_{d\mid p_r^{e_r}}\mu^2(d)/\tau(d)\right)\)

OpenStudy (ikram002p):

n=30 f(30)=f( 1*2*3*5)=1+1/2+1/2+1/2+1/4+1/4+1/4+1/8=27/8 lol xD

imqwerty (imqwerty):

i am getting confused :/ 27/64 27/8 13/4 D:

OpenStudy (ikram002p):

nvm i'm going to sleep goodnight :D

imqwerty (imqwerty):

me too :) goodnight :D

OpenStudy (kainui):

n=30 here we go! \(\sum\limits_{d\mid 30}\mu^2(d)/\tau(d)=\left(\sum\limits_{d\mid 2}\mu^2(d)/\tau(d)\right)\left(\sum\limits_{d\mid 3}\mu^2(d)/\tau(d)\right)\left(\sum\limits_{d\mid 5}\mu^2(d)/\tau(d)\right)\) Now let's write out all the terms of both sides of that equation. This is the statement that this left hand side here: \[\frac{\mu^2(1)}{\tau(1)}+ \frac{\mu^2(2)}{\tau(2)}+\frac{\mu^2(3)}{\tau(3)}+\frac{\mu^2(5)}{\tau(5)}+\frac{\mu^2(6)}{\tau(6)}+ \frac{\mu^2(10)}{\tau(10)}+\frac{\mu^2(15)}{\tau(15)}+\frac{\mu^2(30)}{\tau(30)}\] Is exactly equal to this right hand side here: \[\left(\frac{\mu^2(1)}{\tau(1)}+ \frac{\mu^2(2)}{\tau(2)} \right)\left(\frac{\mu^2(1)}{\tau(1)}+ \frac{\mu^2(3)}{\tau(3)} \right)\left(\frac{\mu^2(1)}{\tau(1)}+ \frac{\mu^2(5)}{\tau(5)} \right)\] Multiply those three binomial terms together and see for yourself that they are indeed exactly equal. :D

OpenStudy (kainui):

@imqwerty trust me on this, and write out the multiplication of that bottom expression, and I think you will understand a lot better.

OpenStudy (ikram002p):

huh lol.

OpenStudy (kainui):

just DO IT! lol

imqwerty (imqwerty):

that is \(\large\ (1+\frac{1}{2})^3=\frac{27}{8}\)

OpenStudy (ikram002p):

which means i win :P i just note the formula was correct both of them but we had error in calculation last night so :P

imqwerty (imqwerty):

but the answer should be \(\large\frac{13}{4}\) right and its 27/8 m getting confused where did we had calculation error ?

imqwerty (imqwerty):

and how is the formula correct when its giving the wrong answer

OpenStudy (ikram002p):

why it should be 13/4 ? its (3/2)^3=27/8 :O

ganeshie8 (ganeshie8):

Nice !

imqwerty (imqwerty):

yes so \(\Large\left( 1+\frac{1}{2} \right)^{\omega(n)}\) is correct :D in your next quetions m having difficulty with the \(\sigma(n)\) part

ganeshie8 (ganeshie8):

While we're at this problem of multiplicative functions, can we prove a more general result

ganeshie8 (ganeshie8):

is it easy to prove that \(\mu(n)\) is multiplicative ?

ganeshie8 (ganeshie8):

prove each of below functions are multiplicative 1) \(\tau(n)\) 2) \(\sigma(n)\) 3) \(\mu(n)\)

ganeshie8 (ganeshie8):

\(f(n)\) is multiplicative if \(f(ab) = f(a)f(b)\) whenever \(\gcd(a,b)=1\)

OpenStudy (ikram002p):

we need to show \(\mu(1)=1\) and show that \(\mu(nm)=\mu(m)\mu(n)\) \(\mu(n)=\begin{cases} 1 & \text{ if } n= 1\\ 0 & \text{ if } p^2|n \\ (-1)^{\omega(n)} & \text{ if } n=p_1p_1...p_{\omega(n)} \end{cases} \)

imqwerty (imqwerty):

do we take cases to prove for \(\mu(n)\)?

ganeshie8 (ganeshie8):

\(\mu(1) = 1\) is by definition we need to prove \(\mu(ab) = \mu(a)\mu(b)\) whenever \(\gcd(a,b)=1\)

ganeshie8 (ganeshie8):

we can use cases if that is easy, but it seems we can do it straight without any cases...

OpenStudy (ikram002p):

so from definition \(\mu(1)=1\) we need to show \(\mu(nm)=\mu(n)\mu(m)\) so we have 3*3 cases if n or m =1 its trivial done. let p^2|n w.l.o \(\mu(mn)=\mu(m).\mu(n)\\ 0=0\) also trivial

ganeshie8 (ganeshie8):

Oh rihgt, cases are required yeah...

OpenStudy (ikram002p):

now we need to consider the last case \(gcd(n,m)=1\rightarrow n=p_1...p_\omega(n) , m=q_1...q_\omega(n) , q_i\neq p_i \forall i\) \(\begin{align*} \mu(mn) &= (-1)^{{\omega(m)}+\omega(n)}\\ &=(-1)^{\omega(m)} . (-1)^{\omega(n)} \\ &=\mu(m).\mu(n) \end{align*} \)

OpenStudy (ikram002p):

next function is tau ?